### v2

# Severi-Brauer Bundle Violating the Hasse Principle

Abstract.
We study Severi–Brauer bundles and their Brauer groups. We answer a question posed by Bjorn Poonen in a more general setting by constructing Severi–Brauer bundles that fails the Hasse principle for rational points, without requiring explicit calculations.
Keywords: Rational points, Brauer Manin obstruction, Severi-Brauer bundle
1. Introduction
1.1. Background
Let $X$ be a smooth projective geometrically integral variety defined over a global field $k$. A central question in arithmetic geometry is whether $X$ has a $k$-point. A necessary condition for the existence of such a point is that $X$ has $k_{v}$-points for all places $v$ of $k$. We know that $X(k_{v})\neq\emptyset$ for almost all places $v$ of $k$ [2, Theorem 7.7.2], which allows us to check whether $X({\mathbb{A}}_{k})\neq\emptyset$ in a finite process. However, this condition is not sufficient for many varieties; it is possible for $X(k_{v})\neq\emptyset$ for all places $v$ of $k$ while $X(k)=\emptyset$. This is referred to as the failure of the Hasse principle.
In 1970, Manin introduced a construction that explains the failure of the Hasse principle for all the varieties known at that time, which is now known as the Brauer–Manin obstruction. By a ‘nice’ variety we mean a smooth projective, geometrically integral variety. Although there are counterexamples to the Hasse principle that cannot be captured by the Brauer–Manin obstruction [6], Colliot-Thélène conjectured that for ‘nice’rationally connected varieties, the Brauer–Manin obstruction is the only obstruction to the Hasse principle. When it is known to be the only obstruction, this gives a finite procedure for checking the existence of rational points.
By the functoriality of the Brauer group, we have a pairing $\langle\cdot,\cdot\rangle:X(\mathbb{A}_{k})\times\text{Br}X\to\mathbb{Q}/\mathbb{Z}$. The Brauer–Manin set is defined as the set of all adelic points orthogonal to $\text{Br}X$ and is denoted by $X({\mathbb{A}}_{k})^{\text{Br}}$. We have $X(k)\subset X({\mathbb{A}}_{k})^{\text{Br}}\subset X({\mathbb{A}}_{k})$. If $X({\mathbb{A}}_{k})$ is nonempty and $X({\mathbb{A}}_{k})^{\text{Br}}$ is empty, we say that $X$ has Brauer-Manin obstruction to the existence of rational points.
We study Brauer–Manin obstruction in the context of Severi–Brauer bundles. As is well known, computing the Brauer group of a general Severi–Brauer bundle is a difficult problem. Since the Brauer group is a birational invariant for smooth projective varieties, it is natural to try to compute the Brauer group of a Severi–Brauer bundle using only its generic fiber. This work is an attempt in that direction. In [1, Prop. 5.1], Bjorn Poonen proved the existence of a Châtelet surface over any number field that violates the Hasse principle. His proof is based on an explicit defining equation of variety and involves a case-by-case analysis of valuations to establish the existence of local points and the Brauer–Manin obstruction. This was later generalised to Châtelet $p$-folds by Anthony Várilly-Alvarado and Bianca Viray in [7]. In the same paper, Poonen asked [1, Open Question (ii)] if there is a proof of [1, Prop. 5.1] which does not require explicit calculation. We answer this question in a more general setting by proving that there exist Severi–Brauer bundles over number fields that violate the Hasse principle, and that the Brauer–Manin obstruction explains this failure, without using explicit calculations.
Theorem 1.1.
Given an odd prime $p$ and a number field $k$ which contains a primitive $p$th root of unity, there exists a Severi-Brauer bundle over $k$ which has Brauer-Manin obstruction to the existence of rational points. Moreover, the obstruction is captured by a single Brauer class of order $p$.
In the literature, most examples for computing the Brauer–Manin obstruction rely on explicit models. In this work, we study the generic fiber of the Severi–Brauer bundle and the associated cyclic algebra to compute its Brauer-Manin obstruction. This allows us to avoid detailed calculations.
For proper scheme $X$, due to the compactness of $X({\mathbb{A}}_{k})$, if $X({\mathbb{A}}_{k})^{\text{Br}}=\emptyset$, then this is captured by a finite subgroup of $\text{Br}\,X$. We also study the exponent of finite group capturing the Brauer Manin obstruction. Since a single Brauer class captures the obstruction in most cases, we investigate whether there is a bound on the order of that class.
Definition 1.2.
We say ‘a finite group $G$ captures Brauer-Manin Obstruction’ if there exists a ‘nice’ scheme $X$ for which $X({\mathbb{A}})^{G}=\emptyset$ but $X({\mathbb{A}})^{H}\neq\emptyset$ for all proper subgroups $H$ of $G$.
Most of the examples of Brauer-Manin obstruction in the literature are captured by a cyclic subgroup of $\text{Br}X$. Recently, Jennifer Berg and their collaborators in an inspiring paper [3], have shown that the subgroup may require arbitrarily many generators. In particular, for any given natural number $n$, $({\mathbb{Z}}/2)^{n}$ captures the obstruction. This motivates us to the following question.
Question 1.3.
Which finite abelian groups can capture the Brauer-Manin Obstruction?
Corollary 1.4.
For any odd prime $p$, the group ${\mathbb{Z}}/p{\mathbb{Z}}$ can capture the Brauer-Manin obstruction. Therefore, the exponent of a finite group capturing Brauer-Manin obstruction has no upper bound.
1.2. Overview of the paper
Throughout the paper, $k$ is a number field that contains a primitive $p$th root of unity for a fixed odd prime $p$, unless stated otherwise. We begin section 2 by recalling some key concepts related to Severi-Brauer varieties and Brauer-Manin obstruction. In Section 3, for any choice of a cyclic algebra over $k(t)$, we construct a Severi-Brauer bundle. Then we ensure that the Brauer group of such a variety is a non-trivial $p$-torsion group by base changing to a Henselian local ring and using the purity theorem. In section 4, we consider the cyclic $p$-algebra $(a,f)$, where $a$ and $f$ are arbitrary choices subject to certain technical conditions. There exists a unique Severi-Brauer bundle over a dense subset of ${\mathbb{P}}^{1}_{k}$ associated with it. By choosing a smooth model, we obtain a nice variety over ${\mathbb{P}}^{1}_{k}$ that has global points. To eliminate global points while preserving local points in the variety, we use a pullback method that first appeared in [4] and was later generalized in [3]. This results in a local-global failure, captured by an element of order $p$, described in section 4. The proofs in section 4 are inspired by the work of Jennifer Berg and their collaborators [3].
1.3. Notation
Throughout the article, a variety is a separated scheme of finite type over a field. By a ‘nice’ scheme we mean a smooth projective geometrically integral scheme. For a global field $k$, we denote the completion at each place $v$ of $k$ by $k_{v}$ and its adele ring by ${\mathbb{A}}_{k}$. For a variety $X$ over $k$, $X^{s}$ denotes the base change of $X$ to the separable closure of $k$.
2. Preliminaries
2.1. Brauer group of a field
(See [5] for details) For any field $k$, let $\text{Br}\,k$ denote its Brauer group. $\text{Br}\,k$ classifies finite central division algebras over $k$ upto similarity. Let $A$ be a central simple algebra over $k$. Then $A$ is said to be split over $k$ if the class of $A$ in $\text{Br}\,k$ is zero. Equivalently, $A$ is split if and only if there exists an integer $m>0$ such that $A\simeq M_{m}(k)$ as $k$-algebras.
Given a central simple algebra $A$ of degree $n$ over an arbitrary field $k$, there exists an associated projective $k$ variety $X$ of dimension $n-1$ known as the Severi-Brauer variety associated with $A$. The existence of a $k$-point in $X$ is characterized as follows.
Lemma 2.1.
Let $A$ be a central simple algebra of degree $n$ over an arbitrary field $k$ and $X$ be the associated Severi-Brauer variety over $k$. The following are equivalent.
  1. (a)
    $X$ has $k$-rational point.
  2. (b)
    $X$ is isomorphic to the projective space ${\mathbb{P}}^{n-1}_{k}$ over $k$.
  3. (c)
    $A$ splits over $k$.
Proof.
The equivalence of statements (a) and (b) is a part of Chátelet’s theorem. The implication $(b)$ if and only if $(c)$ follows from the base-point preserving bijection of central simple algebras of degree $n$ over $k$ and Severi-Brauer varieties of dimension $n-1$ over $k$ as explained in [5]. ∎
Cyclic algebras constitute an important subclass of central simple algebras. These are generalization of quaternion algebras to arbitrary degree. When the base field $k$ contains a primitive $m$-th root of unity $\omega$, a degree $m$ cyclic algebra has a nice presentation. Given any $a,b\in k^{\times}$, there exists a cyclic $m$-algebra denoted by $(a,b)_{\omega}$, which can be described as
$\langle x,y\,|\,x^{m}=a,y^{m}=b,xy=\omega yx\rangle$.
We now present a useful result concerning cyclic algebras, whose proof is given in[5].
Lemma 2.2.
Let $k$ be a field containing a primitive $m$-th root of unity $\omega$ and let $a,b\in k^{\times}$. Then the following statements are equivalent.
  1. (a)
    The cyclic algebra $(a,b)_{\omega}$ is split over $k$.
  2. (b)
    The element $b$ is a norm from the field extension $k(\sqrt[m]{a})$.
  3. (c)
    The element $a$ is a norm from the field extension $k(\sqrt[m]{b})$.
Theorem 2.3 (Amitsur).
Let $X$ be a Severi–Brauer variety defined over a field $k$. Then the kernel of the restriction map $r_{X}:\text{Br}\,k\rightarrow\text{Br}\,\kappa(X)$ is a cyclic group generated by the class of $X$ in $\text{Br}\,k$.
The definition of a Brauer group can be generalized to a scheme using étale cohomology.
2.2. Brauer group of a scheme
For any scheme $X$ over a field $k$, define the Brauer group as $\text{Br}X:=H^{2}_{\text{et}}(X,{\mathbb{G}}_{m})$, and its subgroup of constant classes as $\text{Br}_{0}\,X:=im(\text{Br}\,k\rightarrow\text{Br}X)$. We write $\overline{\text{Br}}\,X$ for the quotient $\text{Br}X/\text{Br}_{0}\,X$. Following are some important properties concerning Brauer group of schemes. For more details see [2].
  1. (a)
    For a field $k$, $\text{Br}(\text{Spec}\,k)=\text{Br}\,k$
  2. (b)
[…]
  3. (c)
    The inclusion of the generic point $i:\text{Spec}\,\kappa(X)\to X$ gives an inclusion of the Brauer group $i^{*}:\text{Br}X\to\text{Br}\kappa(X).$
  4. (d)
    Brauer group is a birational invariant for smooth projective varieties.
For any scheme $X$, let $X^{(1)}$ be the set of codimension 1 points. For $t\in X$, we denote the residue field at $t$ by $\kappa(t)$. If $X$ is integral, the function field of $X$ is denoted by $\kappa(X)$. The following theorem helps us to decide when an element of $\text{Br}\,\kappa(X)$ actually belongs to $\text{Br}X$. For a reference, see [2, Theorem 6.8.3]. This is a version of the Grothendieck’s Purity theorem for the Brauer group.
Theorem 2.4 (Purity theorem).
Let $X$ be a regular integral Noetherian scheme. Then the following sequence
$0\rightarrow\text{Br}X\rightarrow\text{Br}\,\kappa(X)\xrightarrow{\oplus_{x}\partial_{x}}\displaystyle\bigoplus_{x\in X^{(1)}}H^{1}(\kappa(x),{\mathbb{Q}}/{\mathbb{Z}})$
is exact.
In particular, if we choose $X$ to be ${\mathbb{P}}^{1}_{k}$ and a cyclic algebra $(a,g)$ of index $p$ in $\text{Br}\,\kappa({\mathbb{P}}^{1}_{k})$, then for closed points of $t\in{{\mathbb{P}}^{1}_{k}}$, $\partial_{t}(a,g)=a^{v_{t}(g)}$ in $\kappa(t)^{\times}/\kappa(t)^{\times p}\subset H^{1}(\kappa(t),{\mathbb{Q}}/{\mathbb{Z}})$.
Theorem 2.5.
Let $X$ be a regular integral scheme and $U\subset X$ be a dense open subscheme. Then we have an exact sequence
$0\rightarrow\text{Br}X\rightarrow\text{Br}\,U\xrightarrow{\displaystyle}\bigoplus_{D}H^{1}(\kappa(D),{\mathbb{Q}}/{\mathbb{Z}})$
where $D$ ranges over the irreducible divisors of $X$ with support in $X\setminus U$ and $\kappa(D)$ denotes the residue field at the generic point of $D$.
Now we recall some terminologies related to Brauer-Manin obstruction.
2.3. Brauer-Manin obstruction
For a variety $X$ over a number field, one can define the Brauer–Manin pairing $\langle\cdot,\cdot\rangle:X(\mathbb{A}_{k})\times\text{Br}\,X\to\mathbb{Q}/\mathbb{Z}$ by summing the local pairings, defined as $((P_{v}),A)\mapsto\sum_{v}\text{inv}_{v}A(P_{v})$. For $A\in\text{Br}X$, we have the following commutative diagram
${X(k)}$${X({\mathbb{A}}_{k})}$00{0}${\text{Br}\,k}$${\bigoplus_{v}\text{Br}\,k_{v}}$${{\mathbb{Q}}/{\mathbb{Z}}}$00{0}$\scriptstyle{A}$$\scriptstyle{A}$$\scriptstyle{\sum\text{inv}_{v}}$
and we define the subset of adelic points $X({\mathbb{A}}_{k})^{A}:=\{(x_{v})\in X({\mathbb{A}}_{k}):\sum_{v}inv_{v}A(x_{v})=0\}.$ For any subgroup $B$ of $\text{Br}X$ denote $X({\mathbb{A}}_{k})^{B}:=\bigcap_{A\in B}X({\mathbb{A}}_{k})^{A}.$ The above diagram shows that
$X(k)\subset X({\mathbb{A}}_{k})^{\text{Br}}\subset X({\mathbb{A}}_{k}).$
We say that there is Brauer–Manin obstruction to the existence of rational points on $X$ if $X({\mathbb{A}}_{k})$ is non-empty but $X({\mathbb{A}}_{k})^{\text{Br}}$ is empty.
For a regular projective variety $X$, the continuity of Brauer-Manin pairing ensures that if $X$ has Brauer-Manin obstruction then there exists a finite subgroup $B\subset\text{Br}\,X$ that captures the obstruction. Let $B$ be a finite subgroup of $\text{Br}\,X$ and $\hat{B}:=\text{Hom}(B,\mathbb{Q}/\mathbb{Z})$. Then we obtain a set of maps $\phi_{v}:X(k_{v})\to\hat{B}$ for each place $v$, defined by $P_{v}\mapsto(A\mapsto\text{inv}_{v}A(P_{v})).$ These maps sum to give a single map $\phi:X(\mathbb{A}_{k})\to\hat{B}.$ Note that, for all but finitely many places $v$, we have $\phi_{v}(X(k_{v}))=\{0\}$[2, Proposition 8.2.1.], which ensures that the sum is well defined.
3. Brauer groups of Severi Brauer bundles
In this section, we describe the construction of a Severi–Brauer bundle based on the choice of a generic fiber and explore its associated Brauer group. Since the Brauer group is a birational invariant for smooth projective varieties, one would want to compute the Brauer group of a Severi–Brauer bundle using only its generic fiber. This work provides an approach to do so. Additionally, since our construction does not rely on explicit polynomial equations, the techniques presented can be applied to a wide range of varieties.
From now on, $k$ is a number field that contains a primitive $p$th root of unity for a fixed odd prime $p$. Given an element $a\in k^{\times}$ and a polynomial $f(u)\in k[u]$, we describe the construction of a nice variety over $k$. Note that, there is a one-one correspondence between homogenous $F\in k[u_{0},u_{1}]$ whose degrees are a multiple of $p$ and $p$th power-free polynomials $f\in k[u]$ given by
$F(u_{0},u_{1})\mapsto F(u,1)$ and $f(u)\mapsto f\left(\frac{u_{0}}{u_{1}}\right)u_{1}^{p\lceil\deg f/p\rceil}.$
Construction 3.1.
Consider the cyclic algebra $(a,f)\in\text{Br}(k(u))$ of degree $p$. By purity theorem 2.4, there exists an open set U $\subset\mathbb{P}^{1}_{k}$ such that $(a,f)\in\text{Br}\,U$. Since the cyclic algebra $(a,f)$ is unramified at all closed points of ${\mathbb{P}}^{1}_{k}$ except possibly at the irreducible factors of $F$, we can take $U=\mathbb{P}^{1}_{k}\setminus V(F)$ where $F$ is the homogeneous polynomial corresponding to $f$. Let $X_{U}$ represent the family of Severi-Brauer varieties over $U$ associated with the algebra $(a,f)$. Note that for each $q\in U$, the fiber $X_{q}\to\text{Spec}\,\kappa(q)$ is the Severi-Brauer variety associated with the cyclic algebra $(a,f(q))\in\text{Br}\,\kappa(q)$. Using Hironaka’s theorem over $k$, we choose a smooth projective model for $X_{U}\to U$ denoted by $\pi:X_{a,f}\to\mathbb{P}^{1}_{k}$. We like to emphasize that its generic fiber corresponds to the Severi-Brauer variety associated with $(a,f)$, which uniquely determines $\pi:X_{a,f}\to\mathbb{P}^{1}_{k}$ up to birationality. Since a Severi-Brauer variety is geometrically integral rationally connected, we observe that $X_{a,f}$ is a nice rationally connected variety over $k$.
Remark 3.2.
The vertical Brauer group of $X/Y$ is defined as
$\text{Br}_{\text{vert}}(X/Y)=\{A\in\text{Br}(X)\mid j^{*}(A)\in\text{Im}[\text{Br}(\text{Spec}\kappa(\eta))\to\text{Br}(X_{\eta})]\},$
where $j$ denotes the natural inclusion of generic fiber $X_{\eta}\to X.$ For families of Severi-Brauer varieties, the full Brauer group is vertical, that is, $\text{Br}_{\text{vert}}(X)=\text{Br}(X)$ [8, Remark 11.1.12]. In particular for $\pi:X_{a,f}\to\mathbb{P}^{1}_{k}$, $\text{Br}(X_{a,f})=\{\pi^{*}(\beta)\in\text{Br}(X):\beta\in\text{Br}k(t)\}$.
The following theorem examine some properties of the Brauer group of $X_{a,f}$.
Theorem 3.3.
Let $f=f_{1}\dots f_{r}$ with $r\geq 2$, where each $f_{i}$ is a distinct irreducible separable polynomial, and $a$ is not a $p$th power in $k[u]/\langle f_{i}(u)\rangle$ for any $i$. Let $F\in k[u_{0},u_{1}]$ be the homogeneous polynomial corresponding to $f$. Consider the variety $\pi:X\to\mathbb{P}^{1}_{k}$, where $X=X_{a,f}$. We have
  1. (a)
    $\text{Br}\,k\rightarrow\text{Br}X$ is injective.
  2. (b)
    $\overline{\text{Br}}X$ is $p$-torsion.
  3. (c)
    Under the map $\pi^{*}:\text{Br}\,\kappa(\mathbb{P}^{1}_{k})\rightarrow\text{Br}\,\kappa(X)$, $(a,f_{i})$ lands in $\text{Br}X$ for all $i$.
  4. (d)
    $\pi^{*}(a,f_{i})$ is nonzero in $\overline{\text{Br}}\,X$ for all $i$.
Proof.
(a) Let $X_{\theta}$ be the generic fiber of $\pi:X\to\mathbb{P}^{1}_{k}$. We know that the kernel of the map $\text{Br}\,k\rightarrow\text{Br}\,X\hookrightarrow\text{Br}\,\kappa(X)$ is equal to the kernel of the composition $\text{Br}\,k\hookrightarrow\text{Br}\,k(t)\rightarrow\text{Br}\,X_{\theta}\hookrightarrow\text{Br}\,\kappa(X)$. Since the kernel of the map $\text{Br}\,k(t)\to\text{Br}\,X_{\theta}$ is $\langle[X_{\theta}]=[(a,f)]\rangle$ by Theorem 2.3 and $\langle X_{\theta}\rangle\cap\text{Br}\,k=\{0\}$, it follows that the restriction $\text{Br}\,k\to\text{Br}\,X_{\theta}$ is also injective. Thus the claim follows.
(b) This part of the proof follows a similar approach to that in [3, Theorem 6.5.(a)], where the case of conic bundles was addressed. We generalize their proof to Severi-Brauer bundles arising from cyclic algebras and describe the necessary details.
Let $L$ be a field extension of $k$ such that $a\in L^{\times p}$. The scheme $X_{L}$ is birational to $\mathbb{P}^{p-1}_{L}\times\mathbb{P}^{1}_{L}$, which implies $\text{Br}X_{L}\cong\text{Br}\,L$. Now let $L=k(\sqrt[p]{a})$ which is a cyclic Galois extension of $k$ of degree $p$ with Galois group $G$. Using the lower term exact sequence obtained using Hochschild–Serre spectral sequence we have $\overline{\text{Br}}\,X\cong H^{1}_{et}(G,\text{Pic}\,X_{L})$, which is annihilated by $|G|=p$. This shows that $\overline{\text{Br}}\,X$ is $p$-torsion.
(c) Without loss of generality, fix an irreducible factor $f_{i}$ of $f$. We show that $\pi^{*}(a,f_{i})$ lands in $\text{Br}\,X$. Let $R$ be the Henselization of $k[t]$ with respect to $f_{i}$ and $K$ be the fraction field of $R$. Let $\pi_{R}:X_{R}\rightarrow\text{Spec}\,R$ be the structure morphism associated with the base change of our scheme $X$ to $\text{Spec}\,R$ with the generic fiber $\tilde{\pi}_{R}:X_{\eta}\rightarrow\text{Spec}\,K$. By Amitsur’s Theorem 2.3 we have, $\tilde{\pi}_{R}^{*}(a,f)=0$ in $\text{Br}X_{\eta}$. Since $\text{Br}\,\kappa(X_{R})\xhookrightarrow{}\text{Br}\,\kappa(X_{\eta})$, we get $\pi_{R}^{*}(a,f)=0$ in $\text{Br}\,\kappa(X_{R})$.
Let $\hat{f_{i}}=f_{i}/f$, then we have $\pi_{R}^{*}(a,f_{i})=\pi_{R}^{*}(a,f)+\pi_{R}^{*}(a,\hat{f_{i}})=\pi_{R}^{*}(a,\hat{f_{i}})$ in $\text{Br}\,\kappa(X_{R})$. Note that for $j\neq i$, $f_{j}$ is a unit in $R$ implies that $\partial_{f_{i}}(a,\hat{f_{i}})=a^{v_{f_{i}}(\hat{f_{i}})}=1$. That is, $(a,\hat{f_{i}})$ lies in the kernel of the residue map $\partial_{f_{i}}$, which implies that $\pi_{R}^{*}(a,\hat{f_{i}})$ is in $\text{Br}\,X_{R}$ by purity theorem 2.4. Thus we can conclude that $\pi_{R}^{*}(a,f_{i})$ is also in $\text{Br}\,X_{R}$. Let $U=\pi^{-1}(f_{i})$ and $V=\pi_{R}^{-1}(f_{i})$, then by using the purity theorem 2.5, we have the following commutative diagram
00{0}${{\text{Br}X}}$${{\text{Br}\,(X\setminus U)}}$${{\bigoplus}_{D}{H^{1}(\kappa(D),{\mathbb{Q}}/{\mathbb{Z}})}}$00{0}${{\text{Br}X_{R}}}$${{\text{Br}\,(X_{R}\setminus V)}}$${{\bigoplus}_{D^{\prime}}{H^{1}(\kappa(D^{\prime}),{\mathbb{Q}}/{\mathbb{Z}})}.}$
Since the last vertical map is an isomorphism, we get $\pi^{*}(a,f_{i})\in\text{Br}\,X$.
(d) We prove this result by contradiction. Assume $\pi^{*}(a,f_{i})=0$ in $\overline{\text{Br}}\,\kappa(X)$ for some $i$. Let $\pi_{\theta}:X_{\theta}\rightarrow\text{Spec}\,k(t)$ be the generic fiber of $\pi$. By assumption, we have $\pi_{\theta}^{*}(a,f_{i})=0$ in $\overline{\text{Br}}\,\kappa(X_{\theta})$. By Amitsur’s theorem 2.3 and going modulo the image of $\text{Br}\,k$, we obtain the exact sequence
$1\rightarrow\overline{\langle(a,f)\rangle}\rightarrow\overline{\text{Br}}\,\kappa(\mathbb{P}^{1}_{k})\xrightarrow{\pi_{\theta}^{*}}\overline{\text{Br}}\,\kappa(X_{\theta}).$
Since $\pi_{\theta}^{*}(a,f_{i})=0$ in $\overline{\text{Br}}\,\kappa(X_{\theta})$, we have $(a,f_{i})\in\text{Ker}\,\pi_{\theta}^{*}$. This implies $A:=[(a,f^{n})]-[(a,f_{i})]\in\text{Br}\,k$ for some $n$. If $n$ is coprime to $p$ let $j\neq i$ be an element of $\{1,\dots,r\}$, then we have $\partial_{f_{j}}(A)=a^{n}$. As $a$ is not $p$th power in $\kappa(f_{j})$, $a^{n}$ is not a $p$th power in $\kappa(f_{j})$, thus $A\notin\text{Br}\,k$. If $n$ is multiple of $p$ we consider $\partial_{f_{i}}(A)=a^{-1}$. Since $a$ is not $p$th power in $\kappa(f_{i})$, we again have $A\notin\text{Br}\,k$. In both cases, we obtain a contradiction to the implication that $A\in\text{Br}\,k$, which follows from the initial assumption. Hence our initial assumption is wrong and we conclude that $\pi^{*}(a,f_{i})$ is a non-trivial element in $\overline{\text{Br}}\,\kappa(X)$. ∎
Note that the construction of $X_{a,f}$ depends only on the generic fibre. Thus for any given central simple algebra $\alpha$ over an arbitrary number field, one gets a Severi-Brauer bundle over $\mathbb{P}^{1}_{k}$ using Hironaka’s theorem, say $X_{\alpha}$.
Corollary 3.4.
Let $k$ be a number field. Given a cyclic algebra $\alpha$ over $k(t)$ with index $m$, the Brauer group of the associated Severi-Brauer bundle $X_{\alpha}$ is $m$-torsion.
Proof.
Same as the proof of 3.3 (b). ∎
4. Proof of main theorem
In this section, we prove Theorem 1.1 by constructing a Severi-Brauer bundle $X$ for each odd prime $p$. We start with a bundle that has a global point and then modify the image of the map arising from Brauer-Manin pairing at the local level. This is done by pulling back via a suitable polynomial $g(x)\in k[x]$ to ensure that an element of $\text{Br}X$ obstructs the Hasse principle. By using the pullback method, we can modify our variety in a way that allows us to control the images of the map obtained from Brauer-Manin pairing at finitely many places. This, in turn, enables us to control which subgroup of the Brauer group captures the obstruction.
We begin by constructing a variety as in 3.1 by carefully choosing a polynomial $f(x)\in k[x]$ and a constant $a\in k^{\times}$. The choice of the polynomial is guided by Lemma 4.1. Then we can choose the constant depending upon the polynomial. This ensures that the variety $X_{a,f}$ has a $k$-point and a non-zero evaluation map at a non-archimedean place of $k$.
Lemma 4.1.
For any given irreducible polynomial $g(x)\in k[x]$ of degree greater than 3, there exists a constant $c\in k$, a non-archimedean place $v$ of $k$, and a polynomial $h(x)\in k[x]$ such that $g(c)$ is not a $p$-th power in $k_{v}$, but $f(c)=g(c)h(c)$ is a $p$-th power in $k_{v}$, where $f(x)=g(x)h(x)$.
Proof.
Consider the curve $C$ defined by homogenisation of $y^{p}-g(x)$ in ${\mathbb{P}}^{2}_{k}$. Let $d$ denote the degree of $C$. We know that the genus of $C$ is $(d-1)(d-2)/2$ [9, Proposition 2.15], which is greater than 1 by the hypothesis on the degree of $g$. According to Falting’s theorem, $C$ has only finitely many $k$ points. In particular, there exists a $c\in k$ such that $g(c)$ is not the $p$-th power in $k$. Consequently, by the local-global principle, there is a place $v$ of $k$ such that $g(c)$ is not the $p$-th power in $k_{v}$.
Consider $l\in k$ such that $g(c).l$ is a $p$-th power in $k_{v}$ and let $h(x)$ be a polynomial in $k[x]$ such that $h(c)=l$ (we can choose this by interpolation method, for example, $h(x)=x-c+l$). Taking $f(x)=g(x)h(x)$ yields $f(c)=g(c).l$, which is a $p$-th power in $k_{v}$. ∎
We fix an odd prime $p$ and consider a monic irreducible polynomial $f_{1}(x)$ of degree greater than 3. By above lemma, there exist a monic polynomial $h(x)$, a constant $c\in k$, and a non-archimedean place $v$ such that $f_{1}(c)$ is not a $p$-th power in $k_{v}$, but $f(c)$ is a $p$-th power in $k_{v}$, where $f(x)=f_{1}(x)h(x)$. Let $h(x)=f_{2}(x)\cdots f_{n}(x)$, where each $f_{j}(x)$ is an irreducible factor of $h(x)$. Next we choose [^1] Any finite degree extension of $k$ can contain $p$th power of only finitely many elements of the set $\{v\ell:\ell\text{ are primes in }k\}$ an element $a\in k^{\times}$ such that $a$ is not a $p$-th power in $k[u]/\langle f_{i}(u)\rangle$ for each $i\in\{1,2,\dots,n\}$, and $v(a)=1$. For this choice of $a$ and $f$, let $\tilde{X}$ be the variety $X_{a,f}$ over $\mathbb{P}^{1}$ constructed as in 3.1.
In the next proposition, we examine certain local properties of our nice rationally connected variety $X_{a,f}$.
Proposition 4.2.
The variety $\tilde{\pi}:\tilde{X}\to\mathbb{P}_{k}^{1}$ constructed above satisfy the following.
  1. (a)
    For all Archimedean places $w$, the Severi-Brauer bundle $\tilde{X}_{k_{w}}$ is birational to a $\mathbb{P}^{p-1}$ bundle.
  2. (b)
    There exists a non-Archimedean place $v$ of $k$ not lying over $p$ and a specific subgroup $H$ of $\mathrm{Br}\,\tilde{X}$ such that the map $\tilde{\varphi}_{v}:\tilde{X}(k_{v})\to\mathrm{Hom}(H,\mathbb{Q}/\mathbb{Z})$ is nonzero.
  3. (c)
    For all places $\omega$, we have $0\in\,im\tilde{\varphi}_{\omega}$.
Proof.
(a) The generic fiber of $\tilde{X}$ is the Severi-Brauer variety associated with the cyclic algebra $(a,f)$. After base change to an Archimedean place $w$, $a$ becomes a $p$th power, that is the algebra $(a,f)$ splits over $k_{w}$. Using Lemma 2.1, we can conclude that the resulting scheme $\tilde{X}_{k_{w}}$ is birational to a $\mathbb{P}^{p-1}_{k_{w}}$ bundle.
(b) Let $v$ be the place coming from Lemma 4.1. Then $f(c)$ is a $p$th power in $k_{v}$, so the cyclic algebra $(a,f(c))$ splits in $\mathrm{Br}\,k_{v}$, thus the fiber $\tilde{X}_{c}$ has a $k_{v}$-point, say $P$. Let $H$ be the subgroup generated by $(a,f_{1})$ inside $\text{Br}\,\tilde{X}$. Consider the map $\tilde{\varphi}_{v}:\tilde{X}(k_{v})\to\mathrm{Hom}(H,\mathbb{Q}/\mathbb{Z})$ where $\tilde{\varphi}_{v}(P)$ sends $(a,f_{1})$ to $\mathrm{inv}_{v}(a,f_{1}(c))$, as $P$ is above $[c:1]\in{\mathbb{P}}^{1}(k_{v})$ and $P^{*}(a,f_{1})=(a,f_{1}(c))$. Now as $f_{1}(c)$ is not a $p$th power, $k_{v}(\sqrt[p]{f_{1}(c)})$ is a degree-$p$ unramified extension of $k_{v}$, say $L$. Then valuation of the norm group is given by $v(N(L^{\times}))=p\mathbb{Z}$. By our choice of $a$ we have $v(a)=1$, which implies that $a$ does not belong to the norm group. Consequently, by Lemma 2.2, $(a,f_{1}(c))$ does not split in $k_{v}$, thus $\mathrm{inv}_{v}(a,f_{1}(c))$ is non-zero.
(c) Since $f$ is monic, the variety $\tilde{X}$ has a $k$-point $Q$ over infinity which implies $Q^{*}(a,f_{1})=(a,1)=0$. It follows that $\tilde{\varphi}_{w}(Q)=0$ for all places $\omega$. ∎
In the following theorem, we construct a Severi-Brauer bundle that violates the Hasse principle by pulling back $\tilde{X}$ via a polynomial $g$.
Theorem 4.3.
There exists a smooth projective geometrically integral rationally connected variety $X$ over $k$ such that
  1. (a)
    $\overline{\text{Br}}\,X$ is a non-trivial $p$-torsion group,
  2. (b)
    $X$ has $k_{\nu}$ points for all places $\nu$ of $k$,
  3. (c)
    The Brauer-Manin set $X(\mathbb{A}_{k})^{B}$ is empty, where $B$ is a subgroup of $\overline{\text{Br}}\,X$.
Proof.
This proof is motivated from [3, Theorem 7.1]. We apply the techniques from their proof to our context and provide the details for the convenience of the reader. Let $\tilde{X}$ over $\mathbb{P}_{k}^{1}$ be the variety considered in Proposition 4.2. Let $H=\langle(a,f_{1})\rangle\subset\mathrm{Br}\,\tilde{X}$. By Proposition 4.2 (b), we know there exists a non-archimedean place $v$ of $k$ such that the map $\tilde{\varphi}_{v}:\tilde{X}(k_{v})\rightarrow\text{Hom}(H,\mathbb{Q}/\mathbb{Z})$ is non-zero. In particular $\tilde{\varphi}_{v}$ is non-constant. For $1\leq i\leq r$, let $\nu_{i}$ be the non-Archimedean places where the map $\tilde{\varphi}_{\omega}$ is non-constant with $\nu_{1}=v$. Since $0\in im(\tilde{\varphi}_{\omega})$ for all places by Proposition 4.2, we define sets $S_{i}\subset\text{im}(\tilde{\varphi}_{\nu_{i}})$ as, $S_{1}:=\text{im}(\tilde{\varphi}_{\nu_{1}})\setminus\{0\}$ and $S_{i}:=\{0\}$ for $2\leq i\leq r$. Part (a) of Proposition 4.2 ensures that $S_{1}\neq\emptyset$.
For $1\leq i\leq r$, let $\mathcal{U}_{\nu_{i}}:=\tilde{\pi}(\tilde{\varphi}_{\nu_{i}}^{-1}(S_{i}))$ be the non-empty open subset of ${\mathbb{P}}^{1}(k_{\nu_{i}})$ as described in [3, Theorem 7.1]. In the case where $\nu_{1}=v$, we note that $\mathcal{U}_{\nu_{1}}$ is the union of non-empty open subsets $\mathcal{U}_{\nu_{1},s}$ of ${\mathbb{P}}^{1}(k_{\nu_{1}})$ defined by $\mathcal{U}_{\nu_{1},s}:=\tilde{\pi}(\tilde{\varphi}_{\nu_{1}}^{-1}(s))$ for each $s\in S_{1}$. Using weak approximation, we can choose a $k$-point in $\mathbb{P}^{1}$ that lies outside $V(F)$ and $c\in\mathcal{U}_{\nu_{i}}$ for every $i$. We assume that $c=\infty$ after making an appropriate change of coordinates on $\mathbb{P}^{1}$. This ensures that the fiber $\tilde{X}_{\infty}$ is a Severi-Brauer variety, and $\tilde{X}_{\infty}(k_{\nu_{i}})\neq\emptyset$ for all $1\leq i\leq r$. Suppose $\theta_{j}$ are the roots of $f_{j}$ in some finite extension of $k$. For $j=\{1,2,\dots,n\}$, we can choose distinct non-Archimedean places $\omega_{j}$ such that $\omega_{j}\notin\{\nu_{1},\dots,\nu_{r}\}$, $\tilde{X}_{\infty}(k_{\omega_{j}})\neq\emptyset$, and $\omega_{j}$ splits completely in $k(\sqrt[p]{a},\theta_{j})$.
Now consider the following conditions on a polynomial $g(x)\in k[x]$:
  1. (1)
    For each $j$, the polynomial $g-\theta_{j}$ is irreducible and separable over $k_{w_{j}}$;
  2. (2)
    For each $i$, we have $g(P^{1}(k_{v_{i}}))\subset U_{v_{i}}$ and $g(P^{1}(k_{v_{1}}))$ meets $U_{v_{1},t}$ for each $t\in S_{1}$;
  3. (3)
    For each place $w$ such that $\tilde{X}_{\infty}(k_{w})=\emptyset$, the set $g(P^{1}(k_{w}))$ meets $\tilde{\pi}(\tilde{X}(k_{w}))$.
There exist local polynomials satisfying each of the above conditions, in particular existence of local polynomials satisfying conditions (2) and (3) are guaranteed by [3, Lemma 4.3]. Further note that all three conditions above are open in finitely many places, thus using weak approximation we can choose a polynomial $g\in k[x]$ which satisfies them all. Let $\pi:X=X_{a,{f}\circ g}\rightarrow\mathbb{P}^{1}$ be the pullback of $\tilde{X}\rightarrow\mathbb{P}^{1}$ along $g$. By construction of $X_{a,{f}\circ g}$ described in 3.1, we know that $X$ is a nice variety over $k$.
Condition (1) on $g$ and [3, Lemma 4.4] implies that each ${f}_{i}\circ g$ is irreducible and separable over $k(\sqrt[p]{a})$. Consequently, $a$ is not a $p$th power in $k[u]/({f}_{i}\circ g)$. Hence, by Theorem 3.3, we conclude that $\overline{\text{Br}}\,X$ is a non-zero $p$-torsion group. This proves part (a). Condition (2) on $g$ guarantees that $X(k_{\omega})\neq\emptyset$ for each place $\omega$ where $\tilde{X}_{\infty}(k_{\omega})=\emptyset$. Thus we conclude $X(\mathbb{A}_{k})\neq\emptyset$. This proves part (b). Let $\tilde{g}$ be the induced map from $X$ to $\tilde{X}$ corresponding to the polynomial $g$. Note that for all places $\omega$ of $k$, the map $\varphi_{\omega}=\tilde{g}\circ\tilde{\varphi}_{\omega}:X(k_{\omega})\rightarrow\text{Hom}(B,\mathbb{Q}/\mathbb{Z})$ is defined for the subgroup $B=\langle(a,f_{1}\circ g)\rangle$ of $\overline{\text{Br}}\,X$. For each $1\leq i\leq r$, Condition (3) on $g$ ensures that the image of $\varphi_{\nu_{i}}$ is precisely $S_{i}$. For places $\omega\notin\{\nu_{1},\dots,\nu_{r}\}$, we have $\text{im}(\tilde{\varphi}_{\omega})=\{0\}$ which implies that $\text{im}(\varphi_{\omega})=\{0\}$ for all places $\omega$ except $\nu_{1}$. Hence it follows that 00 does not belong to the image of the map $\varphi:X(\mathbb{A}_{k})\rightarrow\text{Hom}({B},\mathbb{Q}/\mathbb{Z})$. In other words, $X(\mathbb{A}_{k})^{B}=\emptyset$. This proves part (c). ∎
Now we are in a position to prove our main theorem.
Proof of Theorem 1.1.
Consider the Severi-Brauer bundle $X$ over $k$ from Theorem 4.3. Since $X(\mathbb{A}_{k})\neq\emptyset$ and $X(\mathbb{A}_{k})^{B}=\emptyset$ where $B=\langle(a,f_{1}\circ g)\rangle$, we conclude that $X$ has Brauer-Manin obstruction to the existence of $k$-rational points. Moreover the obstruction is captured by a single Brauer class of order $p$. Hence the proof. ∎
Proof of Corollary 1.4.
Consider the nice $k$-variety $X$ constructed in Theorem 4.3. The fact that $\overline{\text{Br}}\,X$ is $p$-torsion implies that all of its non-trivial elements have order $p$. In particular, the subgroup $\langle(a,f_{1}\circ g)\rangle$, which captures the obstruction is isomorphic to ${\mathbb{Z}}/p{\mathbb{Z}}$. This concludes that ${\mathbb{Z}}/p{\mathbb{Z}}$ can capture Brauer-Manin obstruction for any odd prime $p$. ∎
Statements and declarations: The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgement: M. Biswas was partially supported by PMRF (ID: 0703045). D.C.R was partially supported by DST-INSPIRE Fellowship (Reg No: IF210208). B. Samanta was partially supported by PhD Fellowship (File No: 09/936(0315)/2021-EMR-I) of CSIR, India. The authors thank the organisers of the ‘Instructional Workshop on Rational Points,’ funded by NWO through an ENW-XL grant, held at Groningen for providing a platform for valuable discussions. B. Samanta is thankful to Infosys travel grant for supporting his participation in the workshop. The authors wish to express their gratitude to Amit Hogadi for his continuous support and guidance throughout this project. The authors are grateful to Anthony Várilly-Alvarado, Brendan Creutz, and Bianca Viray for their valuable comments and suggestions.